### v2

# Taming the Untamed: Graph-Based Knowledge Retrieval and Reasoning for MLLMs to Conquer the Unknown

Abstract
The real value of knowledge lies not just in its accumulation, but in its potential to be harnessed effectively to conquer the unknown. Although recent multimodal large language models (MLLMs) exhibit impressing multimodal capabilities, they often fail in rarely encountered domain-specific tasks due to limited relevant knowledge. To explore this, we adopt visual game cognition as a testbed and select “Monster Hunter: World” as the target to construct a multimodal knowledge graph (MH-MMKG), which incorporates multi-modalities and intricate entity relations. We also design a series of challenging queries based on MH-MMKG to evaluate the models’ ability for complex knowledge retrieval and reasoning. Furthermore, we propose a multi-agent retriever that enables a model to autonomously search relevant knowledge without additional training. Experimental results show that our approach significantly enhances the performance of MLLMs, providing a new perspective on multimodal knowledge-augmented reasoning and laying a solid foundation for future research. [^1] The dataset and code will be released soon.
1 Introduction
Recent multimodal large language models (MLLMs) [54], particularly closed-source ones [1, 2], have demonstrated human-like multimodal capabilities, achieving outstanding performance on benchmarks related to commonsense [42], scientific facts [55], etc. Meanwhile, for domain-specific tasks or rarely seen data [14, 9], relying solely on their own perception and built-in knowledge is inadequate for predicting accurate answers and is hardly interpretable [49, 19].
To alleviate these issues, multimodal RAG (mRAG) [57, 58] has been introduced to improve MLLM performance through heuristic [20, 10, 6] or agent-based methods [44, 33]. However, as with RAG for LLMs [18], such knowledge retrieval approaches often suffer from knowledge redundancy and low relevance [40]. As a result, knowledge graph retrieval [28, 15] has gained increasing attention. Knowledge graphs (KGs) [24], storing knowledge in a structured and comprehensive manner, offer models with more context-rich information. As a substantial multimodal extension, multimodal KGs (MMKGs) [13, 38] can offer a knowledge foundation for enhancing MLLMs.
Figure 1: Our MH-MMKG is curated by knowledgeable players. By leveraging a multi-agents retriever, the MLLM’s responses can be augmented.
One major challenge MLLMs with MMKGs is the unavailability of robust benchmarks. Existing work primarily relies on VQA datasets [29] or web resources [5, 17] for constructing MMKGs. It is highly plausible that such knowledge has already been learned by current MLLMs, making them less useful in evaluating the effectiveness of MMKGs. Additionally, existing MMKGs primarily incorporate text and images, lacking more information-rich modalities, such as video [39], which reduces their suitability for increasingly complex multimodal tasks.
Another challenge is how to retrieve knowledge accurately. The mainstream approach embeds entities in MMKGs to find relevant subgraphs [23, 21, 31, 30]. However, training an effective retriever is data-intensive and unrealistic for low-resource domains. In addition, incorporating graphs about rarely seen domains into powerful closed-source models remains impractical.
This work aims to endeavor the ability of MLLMs to tackle domain-specific tasks by leveraging well-structured external knowledge sources. To achieve this, we build a testbed with a well-known game title [9] as illustrated in Figure 1 for two reasons: First, the visual modality exhibits a large gap from real-world scenes as its visual world is computer graphics-generated mostly with imaginary cultures, creatures, etc. to which MLLMs’ perception module is not well exposed. Second, the knowledge of the game’s world can be different from the real world, which makes the MLLMs’ built-in knowledge less useful. We in this work use “Monster Hunter: World” for our testbed as the game series offers abundant knowledge about the fantasy world. [^2] As one of the most popular game titles, current MLLMs have already learned some knowledge about it. We will experimentally show its impact.
To this end, we construct MH-MMKG, which integrates text, images, videos, and intricate entity relations curated by knowledgeable players. Additionally, we design 238 carefully crafted question-answer pairs as benchmark that cover various sub-tasks, including fine-grained visual cognition, conditional reasoning, etc. Importantly, the knowledge required to answer these questions is encompassed in MH-MMKG. An MLLM can answer most questions with perfect knowledge retrieval and comprehension. MH-MMKG, as well as the benchmark, offers a new dimension of challenges for the community: MH-MMKG requires flexible perception to comprehend the visual modality in an (almost) unseen domain as well as reasoning without relying on the built-in knowledge of the MLLM.
Multi-modal knowledge retrieval through graph is the foundation for tackling this new set of challenges. We thus develop a multi-agent method, which harnesses the self-searching capabilities [26, 45, 16] of MLLMs, allowing them to autonomously retrieve relevant knowledge without training. On both close- and open-source leading MLLMs, we experimentally show the effectiveness of our method, qualifying it as a strong baseline.
Our contribution lies in developing a high-quality benchmark and exploring MLLMs’ ability to find knowledge from MMKG for solving rarely seen domain-specific tasks. These tasks require not only advanced multimodal perception but also a deep understanding of complex dependencies, conditions, and rules within MH-MMKG. Our benchmark, together with the baseline, provides a strong foundation for advancing MLLMs toward real-world challenges.
2 Related Works
2.1 Multimodal Retrieval Augmented Generation
Recent literature shows a growing surge of interest in MLLMs [54]. Despite their advancements, even sophisticated models like GPT-4o face difficulty in handling domain-specific tasks [14, 9] or reasoning [49, 19]. To mitigate these issues, mRAG [57, 58] seeks to enhance AI systems by providing more reliable, comprehensive, accurate, and up-to-date knowledge from external sources [42, 11].
Heuristic mRAG often relies on predefined retrieval strategies that prioritize grounding across multiple modalities into a single primary modality [20, 10, 6, 7], which limits their capacity to deliver precise and contextually rich knowledge. Recent studies [44, 51, 33] propose more adaptable pipelines for knowledge retrieval, utilizing multi-agent cooperation [46] to harness the model’s intrinsic capacity for knowledge exploration. Despite differences in strategy, all these methods depend on retrieving unstructured external knowledge sources, leading to knowledge redundancy and a lack of precision [40]. In contrast, KGs offer a structured format for knowledge storage [24], particularly MMKGs [60], which can be more efficient support for MLLMs to enhance their response.
2.2 Multimodal Knowledge Graph
Compared to KGs, MMKGs incorporate diverse multimodal data, making them more suitable for complex scenarios that require multimodal collaboration. The evolution from conventional KGs to MMKGs has also been extensively explored [13, 38]. Researchers often use densely annotated VQA datasets [29] or web-based resources [5, 17] to automatically construct MMKGs [35, 4]. Ongoing efforts have demonstrated the effectiveness of them in tasks such as VQA [34], image classification [41], cross-modal retrieval [56], etc. Additionally, some studies focused on embedding MMKGs into feature vectors to enhance the reasoning capabilities of MLLMs, yielding impressive results [30, 27].
Notwithstanding these considerable achievements, current works primarily focus on integrating text and image modalities [38] via web data, which is already embedded in the knowledge of MLLMs. The exploration of additional modalities, such as audio and video, as well as practical applications of recent MLLMs for complex real-world tasks, remains limited. Our meticulously crafted MH-MMKG incorporates multiple modalities alongside complex relations as a knowledge base for a domain-specific task.
2.3 Retrieve Knowledge on Graph
By offering reliable and structured information, retrieval over KGs assists LLMs in preserving better factual accuracy [28, 15]. Common approaches include retrieving relevant subgraphs [25, 22] or integrating the graph into the model’s learning process [37]—techniques that have also been extended to MMKGs to improve MLLMs’ task completion capabilities [23, 30]. However, training an effective retriever is highly data-intensive, especially for modalities like video. Moreover, adding new KGs into powerful closed-source MLLMs is infeasible. Recent work aims to evoke the model’s self-searching ability to autonomously navigate toward the necessary knowledge [45, 45, 48]. These approaches typically involve route planning [26, 45] or self-refinement mechanisms [16, 8] to improve retrieval accuracy. We extend these ideas to MMKGs and propose a multi-agent self-searching method for knowledge retrieval.
3 Datasets
The “Monster Hunter” is a popular game series and we choose “Monster Hunter: World” as our testbed to explore the potential of MLLMs in tackling the gaps in visual and knowledge using MMKGs.
3.1 MH-MMKG
Figure 2: The subgraph for monster “Rathian” in MH-MMKG.
Our MH-MMKG is an attribute-based MMKG [60] as shown in Figure 2, where names of elements that appear in the game, such as monsters and attack actions, are treated as entities. A pair of entities can be linked with a particular relation, forming an edge between them. An entity may come with some additional information, i.e., videos related to the entity and textual context to supply more details on the entity, both of which are treated as attributes. Three knowledgeable players were recruited to build the knowledge graph: Each player built subgraphs for single monsters. Figure 2 illustrates the subgraph for “Rathian,” showing intricate relationships such as attack conditions, combos, and inter-subgraph connections. We collected 22 subgraphs, which were merged together to make MH-MMKG.
Let $\mathcal{G}=(\mathcal{E},\mathcal{V},\mathcal{R})$ denote our whole KG, where $\mathcal{E}$ is the set of entities, $\mathcal{V}$ is the set of edges, and $\mathcal{R}$ is the set of relations. $\mathcal{E}$ consists of two types of entities: 1) monsters in $\mathcal{E}_{\text{o}}$ and 2) other entities in $\mathcal{E}_{\text{a}}$. An edge $v=(e,r,e^{\prime})\in\mathcal{V}$ represents $e\in\mathcal{E}$ has relation $r\in\mathcal{R}$ with $e^{\prime}\in\mathcal{E}$. The attribute associated with an entity $e$ is denoted as $A(e)=(c,u)$, which comprises a video clip $c$ and/or textual context $u$. A video clip describes the entity $e$, recorded in 60 fps and 4k. As current MLLMs are not able to handle videos directly, the dataset also offers human-selected keyframes (no more than 10 frames) as well as human-written captions about $c$ [^3] The knowledgeable players who curated the dataset also selected the keyframes and wrote the captions. The textual context $u$ provides knowledge about the entity but is not included in $c$. We denote a set of all attribute as $\mathcal{A}=\{A(e)|e\in\mathcal{E}\}$.
Table 1: The six sub-tasks in our MH benchmark.
Sub-task Description # samples
I: Individual Information Retrieve the textual information of a monster, e.g., nick name, habitat, and skill mechanics. 21
II: Attack Recognition Recognize the about to, ongoing or finished attack action of a monster. 97
III: Combo Premonition Predict upcoming attack sequences based on the monster’s current action or previous action track. 36
IV: Condition Awareness Detect status on monsters or surrounding environment, e.g., whether angry, terrain or phase changes, to anticipate future battle. 30
V: Proc Effect Insight Analyze the effects such as environment and attack on monster status and movement patterns. 40
VI: Cross Monster Analysis Compare attack patterns or behaviors across different monsters to optimize hunting strategies. 14
Figure 3: Our method first converts the query media into a textual caption. Next, a multi-agent self-search mechanism retrieves relevant knowledge associated with the query. Finally, the retrieved knowledge is utilized to enhance the model’s response. Refer to caption means the continue of search and Refer to caption represents the aggregated knowledge upon current entity.
3.2 MH Benchmark
We also designed 238 carefully crafted queries involving knowledge within MH-MMKG. They cover six challenging sub-tasks, as shown in Table 1 (some samples are available in Figure 5). As shown in Figure 3, each input query $Q=(q,d,z)$ consists of 1) a question $q$, 2) a video or images $d$, which serves as a visual reference for $q$, and auxiliary information $z$, which contains relevant monster’s name $e\in\mathcal{E}_{o}$ as well as additional information that cannot be inferred purely from $d$. The auxiliary information $z$ provides additional context to an MLLM for finding the correct subgraph. By default, we assume that $z$ is available as it facilitates knowledge retrieval and in the real scenarios players generally know them [^4] We analyze their impacts in the supplementary material.. Note that, visual data in $Q$ is different from MH-MMKG’s.
To answer $q$ without built-in knowledge about the game’s world, a model needs to go through multiple entities and edges of $\mathcal{G}$. For example, to answer question “Which attack follows?” with a visual context in which a monster Rathian perform attack action Triple Rush, the knowledge will be:
[…]
as shown in Figure 2. Therefore, finding the correct paths over $\mathcal{G}$ is essential for this task. Each query is annotated with a subgraph $\mathcal{I}$ of $\mathcal{G}$ as ground-truth knowledge for answering $q$, textual descriptions $s$ of $d$ written by the knowledgeable game players, and the ground-truth answer $y$. $\mathcal{I}$ comes with a root entity $e_{0}$, and the paths from $e_{0}$ to each leaf is deemed as a bunch of knowledge. The details of MH-MMKG and the MH benchmark are available in the supplementary material.
3.3 Task Definition
We define three task variants over MH benchmark with different levels of supplementary information.
Knowledgeable (Know.) variant simulates a model with sufficient perception for the game’s visual domain and sufficient (built-in) knowledge about the world to answer the given question. This idealized setting assesses how well the model reasons the answer $\hat{y}$ from perfect external knowledge and perception. Together with $Q$ and $\mathcal{G}$, we provide a model with 1) the annotated subgraph $\mathcal{I}$ as well as 2) the textual description $s$ of $Q$’s visual reference $d$, and 3) the set of all human-written caption for $c$ in $\mathcal{G}$ so that it can access to ones associated with entities in $\mathcal{I}$ without visual perception, i.e.:
$\hat{y}=M(Q,s,\mathcal{I},\mathcal{G},\mathcal{A}).$ (2)
Perceptive variant also assumes a model’s sufficient perception but without knowledge to answer the question. This setting evaluates the model’s knowledge retrieval (i.e., to find $\mathcal{I}$) ability and the knowledge comprehension ability. The model is supplied with human-written caption for all $c$ so that it can find textual descriptions associated with entities, and is required to output both the final response $\hat{y}$ and the retrieved $\hat{\mathcal{I}}$:
$\hat{y},\hat{\mathcal{I}}=M(Q,s,\mathcal{G},\mathcal{A}).$ (3)
Unaided is the most challenging variant, where the model relies entirely on its own visual perception (for both $d$ and $c$) to interpret the visual world and retrieve external knowledge. As with the perceptive variant, the output are both $\hat{y}$ and $\hat{\mathcal{I}}$. This variant is formulated as:
$\hat{y},\hat{\mathcal{I}}=M(Q,\mathcal{G}).$ (4)
4 Method
Figure 3 illustrates the overall pipeline of our method to address our task. It is designed as a baseline method for the MH Benchmark and we describe it for the unaided variant, while it can easily adapted to the others. Given an input query $Q$, the visual reference $d\in Q$ is first converted into a textual caption by a perceiver. Then, a multi-agent retriever, consisting of topic selection, validation, and expansion agents, retrieves relevant knowledge for $Q$. Finally, a summarizer generates $\hat{y}$ using retrieved knowledge.
The perceiver $P$ is designed to transform $d$ in $Q$ into textual description. While retaining the original visual data could provide richer information, on-the-fly perception in the visual modality can incur high computational costs; therefore, we choose to convert $d$ into the text as a more efficient representation. The transformed text is demoted as $\hat{s}=P(Q)$. We detail the multi-agent retriever and summarizer in the following sections.
4.1 Multi-agents Retriever
We design a fully automated search algorithm with three agents to find the subgraph of $\mathcal{G}$ to answer $q$. First, the topic selection agent $L$ analyzes the input question $q$ to identify the topic entity $e_{0}=L(q,z,\mathcal{E}_{o})$, which serves as the root for knowledge retrieval. Then, the expansion and validation agents are alternately activated to grow the subgraph with breadth first search, starting from $e_{0}$, to include necessary knowledge. The expansion agent finds all plausible neighboring entities for each open entities. The validation agent, in turn, validate if the current path from the root to each open entities is enough to answer $q$.
We denote the subgraph of $\mathcal{G}$ after the $t$-th round by $\mathcal{K}_{t}=(\mathcal{E}^{\prime}_{t},\mathcal{V}^{\prime}_{t})$, where $\mathcal{E}^{\prime}_{t}\subseteq\mathcal{E}$, $\mathcal{V}^{\prime}_{t}\subseteq\mathcal{V}$, and $\mathcal{K}_{0}=(\{e_{0}\},\emptyset)$. We also denote the set of open entities after the $t$-th round by $\mathcal{O}_{t}\subset\mathcal{E}^{\prime}_{t}$, which is a (sub)set of newly added entities at the $t$-th round and is required to explore further expansion.
For the $(t+1)$-th round, the expansion agent $W$ retrieves the set $\mathcal{N}(e)$ of all plausible neighboring entities for each open entity $e\in\mathcal{O}_{t}$ as:
$\mathcal{N}(e)=W(e,\mathcal{K}_{t};Q,\mathcal{G},\mathcal{A}),$ (5)
where $W$ is an MLLM with a designated prompt (detailed in the supplementary material) to judge if the knowledge aggregated over the path on $\mathcal{G}$ from $e_{0}$ to each $e^{\prime}\in\{e^{\prime}|(e,e^{\prime})\in\mathcal{V}\}$ is useful to answer $q$. $\mathcal{N}(e)$ is a set of entities that are judged to be useful. To build $\mathcal{K}_{t+1}=(\mathcal{E}^{\prime}_{t+1},\mathcal{V}^{\prime}_{t+1})$, we first aggregate all neighboring entities and add them to $\mathcal{E}^{\prime}_{t}$ as:
$\mathcal{E}^{\prime}_{t+1}=\mathcal{E}^{\prime}_{t}\cup\{e^{\prime}|e\in\mathcal{O}_{t},e^{\prime}\in\mathcal{N}(e)\},$ (6)
where duplicated entities are removed to make entities in $\mathcal{E}^{\prime}_{t+1}$ unique. $\mathcal{V}^{\prime}_{t+1}$ includes additional links to newly added entities, given by:
$\mathcal{V}^{\prime}_{t+1}=\mathcal{V}^{\prime}_{t}\cup\{\nu(e,e^{\prime})|e\in\mathcal{O}_{t},e^{\prime}\in\mathcal{N}(e)\},$ (7)
where $\nu(e,e^{\prime})$ gives $v\in\mathcal{V}$ identified by $(e,e^{\prime})$.
The validation agent $U$ checks if each path from $e_{0}$ to $e\in\mathcal{O}_{t+1}$ provides sufficient knowledge to answer $q$:
$o(e)=U(e,\mathcal{K}_{t+1};Q,\mathcal{G},\mathcal{A}),$ (8)
[…]
The open set $\mathcal{O}_{t+1}$ basically is the set of newly added entities in the $(t+1)$-th round. If a newly added entity $e$ is already in $\mathcal{E}^{\prime}_{t}$ and also in $\mathcal{O}_{t}$, the subgraph forms a loop. In this case, $e$ does not need further expansion as it is already in $\mathcal{E}^{\prime}_{t}$. Meanwhile, $e$ may be included in the neighbor sets of multiple $e\in\mathcal{O}_{t}$. In this case, the subgraph also forms a loop, but $e$ has not yet been explored. Thus, it should be in $\mathcal{O}_{t+1}$. Taking these cases into account, $\mathcal{O}_{t+1}$ is given by:
$\mathcal{O}_{t+1}=\mathcal{E}^{\prime}_{t+1}\backslash\mathcal{E}^{\prime}_{t},$ (9)
[…]
This multi-agent retrieval runs until no open entity is found (i.e., $\mathcal{O}_{t}=\emptyset$). The retrieved subgraph is given by:
^$\hat{\mathcal{I}}=\mathcal{K}_{t}.$ (10)
4.2 Reasoning via Knowledge Augmentation
For answer reasoning, we aggregate knowledge on $\hat{\mathcal{I}}$ and represent it in text, which is fed into an LLM. The knowledge augments the reasoning process: the LLM does not need to rely on built-in knowledge about the game’s world but (almost) pure reasoning ability is required. This is formally denoted by:
$\hat{y}=\text{MLLM}(Q,\aleph(\hat{\mathcal{I}},\mathcal{G},\mathcal{A},\alpha)),$ (11)
where $\hat{y}$ is the predicted answer for $Q$, and $\aleph$ transforms each path from $e_{0}$ to each leaf of $\hat{I}$ into text given the entire knowledge (i.e., $\mathcal{G}$ and $\mathcal{A}$). The parameter $\alpha$ limited the number of paths in $\hat{\mathcal{I}}$ used for reasoning (defaulted as 5).
5 Results
Table 2: The experimental settings of query and knowledge retrieval. “Vision” indicates that $d$ is used, while “H-Cap.” uses $s$. For MMKG, “Path” means a model can access annotated subgraphs $\mathcal{I}$. “H-Cap.” means a model use human-written captions (for $c$) in $\mathcal{A}$. “Vis.-Off.” and “Vis.-On.” represent how a model’s captions are generated (offline and online).
Query MMKG
Methods Vision H-Cap. Path H-Cap. Vis.-Off Vis.-On
Vanilla
Vanilla+
Know.
Perceptive
Unaided-Offline
Unaided-Online
Table 3: Experimental results are reported for both leading closed-source and open-source MLLMs. The Vanilla, Vanilla+, and Knowledgeable (Know.) experiments are evaluated solely on Acc., while all other experiments are assessed based on both $Acc.$ and knowledge consistency. Results improved in Online than Offline are highlighted with light green.
Vanilla Vanilla+ Know. Perceptive Unaided-Offline Unaided-Online
Models Acc. Pre. Rec. Acc. Pre. Rec. Acc. Pre. Rec.
GPT-4o [1] .3361 .4033 .9076 .7353 .4545 .6667 .4412 .2500 .4406 .5420 .2715 .5881
GPT-4o mini [1] .3193 .3697 .8361 .6260 .2821 .4952 .4243 .1643 .4434 .3781 .1522 .2835
Claude 3.7 Sonnet [2] .2941 .3403 .9245 .7017 .5681 .6254 .3950 .3002 .3591 .4538 .3069 .4624
Claude 3.5 Sonnet [2] .3151 .3571 .8823 .7142 .5653 .6318 .4160 .3335 .4252 .4075 .2419 .3591
Claude 3.5 Haiku [2] .2605 .3151 .8823 .6344 .3334 .4463 .3781 .1945 .2720 .3907 .1669 .3161
Gemini 2.0 Flash [47] .1890 .2689 .8361 .6596 .3531 .6187 .3823 .1838 .4070 .4159 .1969 .4367
Gemini 1.5 Pro [47] .2394 .2521 .8697 .6932 .4939 .6149 .2899 .1634 .2183 .4075 .2156 .4406
Step-1o [43] .2436 .2815 .8235 .5747 .4372 .5095 .3025 .1831 .2483 .3403 .2204 .2987
InternVL2.5-78B-MPO [12] .1596 .3487 .8949 .6596 .4233 .5076 .3781 .2027 .3007 .3361 .1591 .3235
Qwen2.5-VL-72B [3] .1512 .2731 .8613 .5084 .3284 .3783 .3025 .1622 .1724 .2394 .1463 .1839
Ovis2-16B [36] .1974 .3613 .8655 .5966 .3286 .4406 .2521 .1141 .2817 .2863 .1086 .3122
MiniCPM-o-2.6 [53] .2058 .2857 .8025 .4495 .1761 .2534 .2058 .0590 .1609 .2643 .0594 .2268
DeepSeek-VL2-Small [52] .1344 .2226 .7100 .3058 .0762 .0951 .1890 .0603 .0484 .1512 .0363 .0373
Human (Knowledgeable) .5252 .9033 .9207 .8535
Human (Random) .0336 .6092 .7113 .6457
5.1 Experimental Settings
We evaluated both leading closed-source models (showing in api: gpt-4o-2024-11-20, gpt-4o-mini-2024-07-18 [1], claude-3-7-sonnet-20250219, claude-3-5-sonnet-20240620, claude-3-5-haiku-20241022 [2], gemini-1.5-pro-002, gemini-2.0-flash-001 [47], and step-1o-vision-32k [43]) and open-source models (InternVL2.5-78B-MPO [12], Qwen2.5-VL-72B [3], Ovis2-16B [36], and DeepSeek-VL2 [52]) with our baseline over MH Benchmark. A single model serves as all agents and functionalities (i.e., the perceiver, summarize, as well as the topic selection, validation, and expansion agents in the multi-agent retriever) via different prompts. We use images solely as visual references because current MLLMs do not support videos as input. Due to input size limitations in most models, all images are resized to 1K for our experiments.
[…]
5.2 Performance of MLLMs
In addition to the three tasks defined in 3.3, we evaluate the models’ performance without external knowledge (vanilla). Vanilla+ isolates the effect of the model’s visual ability. Furthermore, the Unaided task includes two variants: 1) Offline variant relies on pre-extracted captions (for all $c$ in MMKG) of an MLLM without access to the query $Q$ or any information obtained during knowledge retrieval. Online variant generates captions for $c$ during knowledge retrieval (i.e., in $U$). Table 2 summarizes the experimental setups. We also include human performance for the vanilla and unaided-online settings. We recruited voluntary participants: One is knowledgeable, and the other has not played the game series. They first answered the questions only with $Q$ and then were allowed to look into $\mathcal{G}$.
As shown in Table 3, with the vanilla setting, all methods hardly predicted correct answers. GPT-4o achieves the best performance with an accuracy of 0.34, whereas the knowledgeable human attains 0.53. The closed-source models generally outperform open-source ones, implying that the closed-source models have richer built-in knowledge about the game. With the vanilla+ setting, which allows human captions $s$ instead of images $d$, most models’ performance improved, which means that MH Benchmark requires comprehension of visual context. In the knowledgeable experiment, the models achieved an accuracy of roughly 0.9. We thus argue that MH-MMKG provides sufficient knowledge to answer the questions in MH Benchmark.
Figure 4: Performance comparison of GPT-4o, Claude 3.7 Sonnet, and Gemini 1.5 Pro across 6 sub-tasks in MH Benchmark.
Figure 5: Examples of the 6 sub-tasks in the MH Benchmark, each generated by GPT-4o for both the Vanilla Answer and the Augmented Answer using an unaided-online retrieval.
The perceptive, unaided-offline, and unaided-online settings evaluate a model’s ability to retrieve relevant knowledge In the perceptive setting, which uses human captions for both queries and knowledge, all models demonstrated a strong capacity to identify relevant knowledge, significantly improving performance compared to vanilla+. Additionally, the models exhibit high knowledge consistency. In the unaided-offline setting, all models performed worse than in the perceptive setting, suggesting that visual perception plays a crucial role in knowledge retrieval. Nevertheless, they are still still better than the vanilla setting, demonstrating the benefits of knowledge retrieval. The unaided-online setting further challenges the models’ visual perception to generate captions with richer context (i.e., $Q$ and the knowledge on the path in $\hat{\mathcal{I}}$) compared to the unaided-offline. Our results show that only some models surpassed the offline variant (highlighted in green), yet the improvement is evident—–GPT-4o improves by 0.1, and Claude 3.7 Sonnet by 0.06. This suggests that knowing what to do enhances a model’s visual perception, strengthening its planning ability, and ultimately results in improved reasoning ability. Additionally, we find that recall is consistently higher than precision, suggesting that models tend to retrieve more knowledge. On the contrary, humans showed higher precision (though both precision and recall are high compared to MLLMs).
We also analyzed the performance differences in the unaided-online setting across six sub-tasks over GPT-4o, Claude 3.7 Sonnet, and Gemini 1.5 Pro, as illustrated in Figure 4. GPT-4o achieves the best performance in all sub-tasks except II. It is particularly strong in sub-tasks IV and V but falls short in sub-tasks II and III, which are more reliant on fine-grained visual perception. Additionally, sub-task I is the simplest, involving only a single relationship $v$, for which the models, except for Gemini, generally perform well. These results suggest that visual perception remains a key challenge for MH benchmark.
Figure 5 presents six examples (one per sub-task) in the vanilla and unaided-online settings, where all predictions are by GPT-4o. In the retrieved path panes, the green paths are both in $\mathcal{I}$ and $\hat{\mathcal{I}}$ (i.e., correctly retrieved paths), the dotted green paths are in $\mathcal{I}$ but not in $\hat{\mathcal{I}}$, and the black paths are in $\hat{\mathcal{I}}$ but not in $\mathcal{I}$. The key relationship or essential information for deriving the correct answer is highlighted in orange. We report the precision and recall metrics for each example. GPT-4o retrieves many knowledge paths, capturing all relevant knowledge for prediction, though precision is low. In Figure 5, we can see that GPT-4o actually retrieved many paths for sub-tasks III and VI. Interestingly, for challenging cases like sub-tasks II and IV, GPT-4o successfully recognized detailed visual concepts (e.g., the monster flying backward and the monster sticking to the wall) and generated plausible captions, which require fine-grained visual perception.
5.3 Analysis of Factors Affecting Performance
Impact of Captioning Performance. Table 4 compares performance scores of captioning, where the metric similarity (Sim.) measures the similarity between generated and human captions using GPT-4o (see supplementary materials for details). The table also summarizes the corresponding reasoning accuracy and knowledge consistency scores. The results indicate that the unaided-online setting consistently gives higher similarity scores than the unaided-offline, suggesting that awareness of query $Q$ and retrieval process consistently enhances captioning performance. Also, the similarity scores positively correlated with the reasoning accuracy scores. We also evaluated the unaided-offline variant on GPT-4o but its offline captions are replaced with ones by video models InternVideo2.5 [50] and VideoChat-Flash [32]. The results showed lower performance than the original GPT-4o across all metrics, indicating that current video models have limited capability in visual comprehension for MH Benchmark.
Table 4: Reasoning and knowledge consistency performance and captioning performance. Improved performance scores due to online captioning are highlighted in green.
Model Vis.-Off Vis.-On Acc. Pre. Rec. Sim.
GPT-4o [1] .4412 .2500 .4406 .2715
GPT-4o [1] .5462 .2715 5881 .2913
Claude 3.7 Sonnet [2] .3950 .3002 .3591 .2649
Claude 3.7 Sonnet [2] .4538 .3069 .4624 .3162
Gemini 1.5 Pro [47] .2899 .1634 .2183 .1643
Gemini 1.5 Pro [47] .4075 .2156 .4406 .1712
InternVideo2.5 [50] .3697 .1960 .2959 .1669
VideoChat-Flash [32] .3445 .2135 .2863 .1421
Keyframe Selection. Due to the limited number of input tokens, current MLLMs cannot handle videos, and keyframe sampling is necessary. MH-MMKG provides human-selected keyframes to facilitate the evaluation on this point. To show the difference between human-selected keyframes and typically-adopted equal-interval sampling, we compare them with GPT-4o and Claude 3.7. Table 5 summarizes the results, where sampling is done in two frames per second, and the maximum number of frames is capped at ten frames. The keyframe selection strategy does impact the performance: Human-selected keyframes seem to provide more informative visual cues, though the difference is not substantial.
Table 5: Impact of Keyframes selection.
Unaided-Offline Unaided-Online
Settings Model Acc. Pre. Rec. Acc. Pre. Rec.
Human GPT-4o .4412 .2500 .4406 .5462 .2715 .5881
Sampling GPT-4o .4117 .2363 .4061 .5037 .2494 .5368
Human Claude 3.7 .3950 .3002 .3591 .4538 .3069 .4624
Sampling Claude 3.7 .3571 .2798 .3248 .4033 .2739 .3749
(a) Using K paths as knowledge.
(b) BFS vs DFS
Figure 6: Ablation experiments for proposed multi-agents search.
The Number of Paths Used in Reasoning. For the same reason, the number of the paths in $\hat{\mathcal{I}}$ used for reasoning is limited. All evaluations so far used 5 random paths, [^6] Our baseline uses the first 5 paths found in the retrieval. A shorter path is thus preferred. though this number can change the performance. Figure 6(a) shows the relationship between the number of paths and the reasoning accuracy. GPT-4o gives the optimal performance when 5 paths are used. We also show in the figure the number of queries $Q$ for which at least a certain number of paths are retrieved. GPT-4o tends to find more paths as its knowledge (the recall is higher), while Claude is more conservative in retrieval (the precision is higher).
BFS versus DFS. The retrieval strategy can also impact performance. BFS finds shorter paths first, while depth-first search (DFS) can find longer paths at the early stage of retrieval. We evaluated the performance for BFS (the proposed baseline) and DFS when three paths are used. As shown in Figure 6(b), BFS consistently performs better. This means that our queries can be answered within fewer hops in $\mathcal{G}$, which can be seen as a limitation of our MH Benchmark, as it only requires fewer steps of reasoning.
6 Conclusion
This work explores the ability of MLLMs to handle domain-specific tasks by retrieving knowledge from MMKG. We introduce MH-MMKG and the MH benchmark as a testbed. Additionally, we propose a baseline with a multi-agent knowledge retriever that allows MLLMs to autonomously access relevant knowledge without requiring additional training. Experimental results on both highlight the importance of visual perception of MLLMs and finding relevant knowledge for reasoning. Our work paves the way for more adaptable and context-aware MLLMs in complex real-world scenarios. Future research may focus on expanding the benchmark size, incorporating a wider range of knowledge types, exploring the potential of video modalities, and developing more advanced knowledge retrieval methods.
Acknowledgement
This work is supported by JST ACT-X Grant Number JPMJAX24C8, JSPS KAKENHI No. 24K20795, No. JP23H00497, CREST Grant No. JPMJCR20D3, JST FOREST Grant No. JPMJFR216O,